The difference of three times the third and two times the second odd integer is 20 less than five times the first odd integer if x is the 1st odd integer. What are the three odd integers?


Answer by badet32(1)

let x=the first odd integer
x+2=second odd integer
x+4=third odd integer
Equation: 3(x+4)- 2(x+2)= 5x - 20
3x+12-2x-4=5x-20
x+8=5x-20
8+20=4x
x=28/4
x=7 first consecutive odd integer
therefore, x+2=7+2=9 second consecutive integer and
x+4=7+4=11 third consecutive integer.
Ans. 7, 9, 11.